OpenStudy (anonymous):

An object in freefall near earths surface accelerates vertically downward at 9.8 m/s due to gravity. This acceleration is also called 1 g. A. If the object falls for 100 m, how fast is It traveling B. How much time is required for it to fall this 100m

OpenStudy (anonymous):

Do you know the expression for position?

OpenStudy (anonymous):

A. v^2-vo^2 = 2gh B.h=1/2g*t^2

OpenStudy (anonymous):

Actually a better expression for A would be v=at+vo

OpenStudy (ghazi):

if it is a case of free fall then your final velocity at any instant would be given by \[V _{f}=V _{0}+g*t\] where Vf is final velocity V0 is initial and t is the time rest g has the same meaning

OpenStudy (ghazi):

by the way in free fall your initial velocity will be equal to 0

OpenStudy (anonymous):

He does not mention the initial velocity, you cant assume that

OpenStudy (ghazi):

he said freefall, that means initial velocity is zero because object has been dropped

OpenStudy (anonymous):

I dont know if thats what freefall means, but even if it is, he could be looking at an instant after the ball was dropped.

OpenStudy (ghazi):

he said that at the height of 100m which means H=0.5*9.8*t^2\[t= \sqrt {\frac{ 100 }{ 9.8*0.5}}\] and then to find Velocity \[V _{f}=9.8 * \sqrt{\frac{ 100 }{ 9.8*0.5 }}\] hope that helps

OpenStudy (ghazi):

\[H=\frac{ 1 }{ 2 }*g*t^2\] for initial velocity = 0m/s

OpenStudy (anonymous):

Oh, I see what you've messed up, the guy who said that was not the one who asked the question, his answer is wrong.

OpenStudy (ghazi):

well, sorry i didn't get ya, but i believe my answer is correct and it will help the asker

OpenStudy (anonymous):

What I said is that the guy who wrote that 0.5Mgt^2=h (that is were you got the vo=0 from), was not the asker, but yes, it will help.

OpenStudy (ghazi):

yea it is me who wrote H= 0.5 *g*t^2